OpenStudy (anonymous):

Greg jogged 3 1/2 miles on monday and 4 3/10 miles on tuesday. How many miles did greg jog altogehter in two days? a. 7 1/3 b. 7 3/4 c. 7 4/5 d. 7 9/10 Can you please look my question

OpenStudy (anonymous):

3 integers + 1/2 + 4 integers + 3/10 =30/10 + 5/10 + 40/10 + 3/10 = 78/10

OpenStudy (anonymous):

The answer above is correct. 7.8 miles or 7 8/10 = 7 4/5.

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

What's the problem O.o

OpenStudy (anonymous):

greg jogged

OpenStudy (anonymous):

The answer is C.